OpenStudy (anonymous):

i^11√-121 Question options: 11 -11 11i -11i

OpenStudy (anonymous):

Simplify

OpenStudy (jdoe0001):

bear in mind that \(11^2 = ?\)

OpenStudy (anonymous):

121

OpenStudy (jdoe0001):

yeap thus \(\bf i^{11}\sqrt{-121}\qquad {\color{brown}{ 11^2\to 121 }}\qquad i^{11}\sqrt{-11^2}\implies i^{11}\sqrt{11^2\cdot -1} \\ \quad \\ i^{11}\sqrt{11^2}\cdot \sqrt{-1}\implies i^{11}\cdot 11\cdot i\implies ?\)

OpenStudy (anonymous):

so its -11?

OpenStudy (jdoe0001):

well... .what would the above give you though.... keep in mind that \(\bf i \implies i^1\)

OpenStudy (jdoe0001):

\(\bf i^{11}\sqrt{11^2}\cdot \sqrt{-1}\implies i^{11}\cdot 11\cdot i^1\implies ?\)

OpenStudy (anonymous):

I really do not know

OpenStudy (anonymous):

can you give me a hint

OpenStudy (jdoe0001):

well.. recall your exponents rules... thus \(\bf i^{11} \cdot i^1 \implies ?\)

OpenStudy (anonymous):

11i

OpenStudy (jdoe0001):

well... what would this \(\bf i^{11}\cdot i^1\implies ?\) give you though?

OpenStudy (anonymous):

1

OpenStudy (jdoe0001):

wel.... what's the exponent of "i" though? when multiplied

OpenStudy (anonymous):

i^2

OpenStudy (anonymous):

x

OpenStudy (jdoe0001):

well... you have an 11 and a 1 exponent recall your exponent rules....same base, different exponents

OpenStudy (jdoe0001):

http://www.math-play.com/image-exponents-rules.jpg

OpenStudy (anonymous):

Can i please know the answer

OpenStudy (anonymous):

i am really confused

OpenStudy (anonymous):

@midhun.madhu1987